OpenStudy (anonymous):

PLEASE HELP !!! --------------------------------------- ~MEDAL WILL BE REWARDED~ --------------------------------------- Simplify the following expression. (53x + 79y) - (-30x + 7y) A. 23x + 72y B. 83x + 72y C. 23x + 86y D. 83x + 86y ---------------------------------------

OpenStudy (anonymous):

(53x + 79y) - (-30x + 7y) We first need to distribute the minus sign. 53x + 79y + 30x - 7y -(-30x) = 30x -(7y) = -7y Now all we need to do is group like terms. 53x + 79y + 30x - 7y 53x + 30x + 79y - 7y 53x + 30x = 83x 79y - 7y = 72y 83x + 72y <----- Answer. (:

OpenStudy (anonymous):

I have some more...

OpenStudy (anonymous):

Simplify the following expression. (50x - 23y) + (38x + 61y) 27x + 84y 27x + 38y 88x + 38y 88x + 84y